Wheat is produced under perfectly competitive conditions. Individual wheat farmers have U-shaped, long-run average cost curves that reach a minimum average cost of $3 per bushel when 1,000 bushels are produced.

b) Suppose demand shifts outward to QD = 3,200,000 - 200,000P. If farmers cannot adjust their output in the short run, what will market price be with this new demand curve? What will the profits of the typical farm be?
1
Expert's answer
2019-07-16T09:31:27-0400

In the long-run free-entry equlibrium p=$3

Demand Q= 3,200,000- (200,000X3)= $2,600,000

Suppose Q=2,600,000, the price=6

Hence firm's make profits= (6X1000)- (3X1000)= $3000